OpenStudy (anonymous):

Q1: Bobby is testing the effectiveness of a new cough medication. There are 100 people with a cough in the study. Seventy patients received the cough medication, and 30 other patients did not receive treatment. Thirty-four of the patients who received the medication reported no cough at the end of the study. Twenty of the patients who did not receive medication reported no cough at the end of the study. What is the probability that a patient chosen at random from this study took the medication, given that they reported no cough? @zepdrix

zepdrix (zepdrix):

Hmm I'm not very good with probability stuff :( Thinking...

OpenStudy (anonymous):

I'm pretty good at guessing, if you can't figure it out :P would you like the options? Maybe that'll help ya out!

zepdrix (zepdrix):

ya that might be a good idea.

OpenStudy (anonymous):

63% 53% 38% 43%

zepdrix (zepdrix):

So when the study was finished, 34/70 reported no more cough (70 being the group that was medicated). 20/30 reported no more cough (30 being the group that received no medical help). So all together, the group of people reporting no more cough is:\[\Large\rm \frac{54}{100}\]54 out of the total 100 people, yes?

OpenStudy (mathmate):

@alg2help Try a contingency table of all the situations.

OpenStudy (anonymous):

Gotcha!

zepdrix (zepdrix):

And then the odds that they were in the medicated group is 70% (70 out of 100).\[\Large\rm \frac{70}{100}\cdot \frac{54}{100}=37.8\text%\]I dont think I did that correctly lol. But uhhhhh whatever

zepdrix (zepdrix):

Yah do some table math.. or whatever mathmate said :d

OpenStudy (anonymous):

Rounded up to 38%, which is what I originally guessed, but cool!!! Thank you to you both! You've been so awesome!!!

OpenStudy (mathmate):



OpenStudy (mathmate):

What is the probability that a patient chosen at random from this study took the medication, \(\text {given that they reported no cough}\)